OpenStudy (anonymous):

if \[x+\frac{1}x=6\]then \[\frac{1}{x^2}+x^2=?\]

OpenStudy (anonymous):

is it correct : \[(x+1/x)^2=6^2\] \[x^2+1/x^2+2=36\] \[x^2+1/x^2=34\]

OpenStudy (anonymous):

m i right ?

hero (hero):

Looks right

OpenStudy (anonymous):

k thanks hero

hero (hero):

But I would have wrote it like this: 1/x^2 + x^2 = 36 and just left it like that

OpenStudy (anonymous):

k

OpenStudy (anonymous):

2 not 2/x?

OpenStudy (anonymous):

\[x^2+\frac{2}{x}+\frac{1}{x^2}\]